Indicate in standard form the equation of the line passing through the given points.
R(3, 3), S(-6, -6)

Answer by Timnewman(323)

Hi,
Since the line passes through these points (3,3) and (-6,-6)
then, the equation of the line will be given by;
(y-y1)(x2-x1)=(x-x1)(y2-y1)
From your question,
y1=3,y2=-6,x1=3,x2=-6
Substitute for y1,y2,x1,x2 in the above formula,i have
(y-3)(-6-3)=(x-3)(-6-3)
(y-3)(-9)=(x-3)(-9)
Clear the bracket
-9y+27=-9x+27
-9y=-9x+27-27
-9y=-9x
Dividing bothsides by -9,
y=x --{This is the required equation}